Read the following information and answer the questions:

-The relationship between the weight of the DNA and the number of nucleotides is: 1 gram of nucleic acid has 2.0 x1021

-The relationship between the pairs of nucleotides and the length of the DNA double strand is: 2.9x106 pairs of nucleotides equal 1mm(millimeter) of DNA double stand.

-The molecular weight of DNA in a haploid cell is called C Value. In humans, the C Value is 3.2x10-12

- An E. Coli cell has 4.7x10-15 grams of DNA double stand on its genome

A) Determine the number of nucleotides that have the C Value in humans ____________________________.

B) The number of paired bases in the human DNA according to what you determined in the premise (A) is: ________________________.

C) Whats going to be the length of the DNA of a haploid human cell? ______________ and the diploid cell?_________________.

D) Determine the number of nucleotides E.Coli has. ______________________

E) Identify the number of nucleotide pairs in its DNA ______________________

F) Whats the size of the DNA molecule in E. Coli? ________________________
Answer #1

A. The number of nucleotide that have the C value in humans

Humans C value is 3.2*10-12 and relation between nucleotide number and molecular weight is 1 gram DNA=2.0*1021

So that no. of nucleotide of human C value is 3.2*10-12*2.0*1021=6.6*109 nucleotides.

  1. The number of paired bases in the human DNA according to what you determined in the premise (A) is:6.6*109/2=3.3*109 in pair no. of bases become half.
  2. The length of human haploid will be1.13*103mm (the relation between length and weight is 2.9*106 base pair is 1mm so that the no. of base pair in haploid is 3.3*109…(A) so the length will be 3.3*109/2.9*106=1.13*103mm) and in diploid cell 1.13*103mm*2=2.26*103mm (because diploid has double set of chromosome)
  3. Number of nucleotides in E coli is 2.0*1021*4.7*10*-15=9.4*106( the relation between number of nucleotides and weight is 1 gram of nucleic acid = 2.0*1021
  4. The number of nucleotide pair in its DNA is9.4*106/2=4.7*106( no. of base pair is half of total bases)
  5. The size of DNA molecule in E . coli is 4.7MB